OpenStudy (anonymous):

without graphing, decide whether the following system of linear equations has one solution, infinitely many solutions, or no solution. 8x=2y-16 and y=4x Explain. Would it not be No solution?

OpenStudy (wilsondanielle):

When you rearrange the first equation to isolate for y, it yields y=4x-16. You are looking for any values of x that mean 4x-16=4x To find solutions, move the terms to the same side and solve for x. 4x-4x = 16 Since 0 is not equal to 16, there are no solutions.

OpenStudy (anonymous):

and if you simplify 8x=2y-16 wouldn't it be y=4x the same as the second part?

OpenStudy (campbell_st):

the simple thing to do is rewrite the 1st equation in slope intercept form 2y = 8x - 16 divide by 2 and y = 4x - 8 now what do you notice about this line and y = 4x

OpenStudy (anonymous):

Yeah! its the same as the second part! :)

OpenStudy (campbell_st):

well almost... both y = 4x - 8 and y = 4x have the same slope... and you should know lines with the same slope don't intersect... so there is no point of intersection, so no solution

OpenStudy (anonymous):

Thank you @campbell_st